A deer is moving with its position varying as a function of time according to the equation: x = 18 + 1.75t – 0.04t 2 Where x is in cm and t is in seconds.

i. What is the deer’s: - initial position? - initial velocity? - initial acceleration?

ii. Determine the time when the deer’s velocity is zero?

iii. How long after starting does the deer’s displacement become zero? Where is the deer then?


1
Expert's answer
2021-05-16T17:59:11-0400

i. "x_0=18\ (cm)" ; "v_0=x'(0)=1.75-0.08\cdot0=1.75\ (cm/c)"

"a_0=v'(0)=-0.08\ (cm/c^2)"


ii. "v=1.75-0.08t\to 0=1.75-0.08t\to t=21.875\ (s)"


iii. "\Delta x=0" if "x=18\ (cm)" . So,


"18=18+1.75t-0.04t^2\to t=43.75\ (s)"

The deer is at the point "x = 18\ (cm)" .